What mass of oxygen is needed for the complete combustion of 6.00×10−3 g

g of methane?


1
Expert's answer
2021-10-20T02:52:12-0400

CH4 + 2O2 → CO2 + 2H2O

M(CH4) = 16 g/mol

n(CH4) "= \frac{6.00 \times 10^{-3}}{16} = 0.375 \times 10^{-3} \; mol"

According to the reaction:

n(O2) = 2n(CH4) "= 2 \times 0.375 \times 10^{-3} = 0.75 \times 10^{-3} \;mol"

M(O2) = 32 g/mol

m(O2) "= 0.75 \times 10^{-3} \times 32 = 24 \times 10^{-3} = 0.024 \;g"

Answer: 0.024 g